Light travels at 2.999 x 108 m/sec in a vacuum. How long will it take (hours) for light to travel from the sun to earth, a distance of 92,955,807 miles?
1
Expert's answer
2016-01-09T12:54:29-0500
Solution
l = 92,955,807 miles
1 mile = 1,609.344 metres
l = 92955807 * 1609.344 = 149597870260.608 (metres)
t = l/v = 149597870260.608 / 2.999*108 = 498.8 (sec) / 3600 = 0.1385 hour

Answer
498.8 sec or 0.1385 hour